# Data Needs and Challenges of Quantum Dot Devices Automation: Workshop Report

### v1

@@ Abstract
Gate-defined quantum dots are a promising candidate system to realize scalable, coupled qubit systems and serve as a fundamental building block for quantum computers. However, present-day quantum dot devices suffer from imperfections that must be accounted for, which hinders the characterization, tuning, and operation process. Moreover, with an increasing number of quantum dot qubits, the relevant parameter space grows sufficiently to make heuristic control infeasible. Thus, it is imperative that reliable and scalable autonomous tuning approaches are developed. In this report, we outline current challenges in automating quantum dot device tuning and operation with a particular focus on datasets, benchmarking, and standardization. We also present ideas put forward by the quantum dot community on how to overcome them.
@@ I Background and motivation
Gate-defined semiconductor quantum dots are a candidate system to realize scalable, coupled qubit systems that serve as a fundamental building block for quantum computers [1, 2]. Their potential for leveraging the materials science and fabrication techniques of the semiconductor industry, while promising remains hard to realize at scale. Specifically, current quantum dot devices suffer from variations in the device dimensions and properties, as well as defects. The corresponding disorder in the electronic landscape throughout a device can be overcome by using local gate voltages to tune the device into an operation regime for quantum computing. However, even tuning a double quantum dot device constitutes a nontrivial task, with each dot typically being controlled by at least three metallic gates, each of which influences the number of electrons in the dot, the tunnel coupling to the outer leads, and the interdot tunnel coupling, which are critical parameters for qubit operation.
@@
The current practice of characterizing and tuning quantum dots for qubit operation either manually or using script-based methods is a relatively time-consuming procedure that is inherently impractical for scaling up and applications. Moreover, with an increasing number of quantum dot qubits, the relevant parameter space grows sufficiently to make heuristic control unfeasible. To overcome the limitations of human-driven experimental control, researchers working with semiconducting quantum dot devices have put considerable effort into automating device control and characterization [3].
@@
To date, several automated tuning algorithms for single- and double-quantum-dot devices have been proposed and demonstrated. While initially all efforts focused on developing in-house script-based algorithms that were tailored to a particular device, more recently members of the community have begun to take advantage of the data analysis tools provided by the field of artificial intelligence and, more specifically, supervised and unsupervised machine learning [3]. When provided with proper training data, machine-learning-driven and machine-learning-enhanced methods have the flexibility of being applicable to various devices with minimal to no adjustments or retraining. Moreover, by learning the governing rules and dynamics directly from the data, such algorithms may be less susceptible to programming errors.
@@
However, machine learning models typically require large, labeled datasets for training, validation, and benchmarking and often lack information about the reliability of the output. Moreover, since the application of machine learning to quantum dot tuning, characterization, and control is a relatively new field of research, it lacks standardized measures of success. The reported success rates vary significantly in both the level and meaning of the reported performance statistics, making it hard (if not impossible) to benchmark the proposed techniques against more traditional tuning approaches or against one another [3]. The time is thus ripe to discuss the broadly defined needs and potential next steps in the field of quantum dot device automation that would enable substantive progress.
@@
When it comes to datasets, most characterization and tuning efforts undertaken to date rely on datasets that consist of either simulated data (which may lack important features representing real-world noise and imperfections) or manually labeled experimental data (which might be subject to qualitative and/or erroneous classification). Moreover, with a few exceptions, experimental datasets are not made publicly available. At the same time, systematic benchmarking of tuning methods on standardized datasets, analogous to the MNIST [4] or CIFAR [5] datasets, is a crucial next step on the path to developing reliable and scalable autotuning algorithms for quantum dot devices.
The goal of the Workshop on Advances in Automation of Quantum Dot Devices Control was to serve as a starting point for discussions about the community’s needs and interests [6]. For two days stakeholders from industry, academia, and the government interested in research and development of semiconductor quantum computing technologies discussed methods of collaboration and future roadmap development of methods for tuning large-scale devices. Topics discussed during the meeting included:
@@
The workshop discussions, guided by a set of questions first put forward in a Federal Register Notice published on April 13, 2023 [7], revolved around five main themes which we present in the following sections. In terms of datasets and benchmarking, the workshop participants were asked to identify public or restricted-use datasets related to the various phases of tuning semiconductor quantum dot devices that are available for training and benchmarking new artificial intelligence models or to test hypotheses using data mining/machine learning methods. They were asked to consider the work researchers need to do to access, and then explore the quality of, existing datasets before conducting research with them and to identify what aspects of this work could be reduced or conducted just once so that future researchers can reduce the time needed to complete a research project. The workshop participants were also asked to consider best practices for creating new datasets or linking existing datasets and sharing them with researchers while adhering to local, State, and Federal laws, as well as barriers and limitations that currently exist to data sharing. Finally, they were asked to think about what role the National Institute of Standards and Technology (NIST) can play in developing infrastructure that supports the use of large-scale datasets for research on tuning quantum dot devices.
@@
In addition to the overall comprehensive summary, each thematic section of this report concludes with a set of core recommendations for possible next steps for advancing automation of the quantum dot tuning and characterization. These recommendations do not represent an endorsement by or an official position of the National Institute of Standards and Technology (NIST) or any other agency participating in the workshop. Rather, they should be understood as summarizing the plurality of views from the workshop.
@@
      NIST could work with the community to define the ontology and core guidelines under the FAIR principles, along with a living process for improving and revising these guidelines, to enable data sharing across the community.
@@ II.1 Science and process challenges
At a high level, quantum dot systems have a variety of engineering and science challenges. However, the largest, pressing challenge is developing a reliable path towards large arrays, in one or two dimensions (or other geometries with scaling at least proportional to the linear size of the system), of quantum dots that are tunable into the quantum coherent, few-electron regime [8].
An example of the challenge ahead was characterized by the difficulties of direct emulation of the Hubbard model. In 2019, Vandersypen and co-authors demonstrated Nagaoka ferromagnetism in a small quantum dot array with four dots [9]. This required building a Hamiltonian-level understanding of the device and maintaining the device in a narrow regime of Hamiltonian parameters throughout the experiment. The time to tune up, and retune, the device becomes a potential limiting factor for scaling to larger numbers of qubits (or, in this case, Hubbard model sites).
@@ II.2 Guidelines for sharing data
Advances in the automation of the characterization and tuning of quantum dot devices cannot happen without the joint effort of the whole community. The development of specialized automation techniques, both machine-learning- and non-machine-learning-driven, is at present stymied by the lack of high-quality relevant datasets. Although some groups already share their data using general-purpose open repositories such as Zenodo [10] or Open Science Framework [11], such data is typically unlabeled and limited in scope to include only the good data that was measured to create the final version of the paper. While a lot of things can go wrong with measurement, the negative counter-examples representing bad data (e.g., poor quality measurement, measurements over incorrectly set parameters) or failure modes are rarely made available. This makes it challenging to learn from mistakes and results in multiple groups reproducing the same errors. Moreover, in some groups, it is still a fairly common practice for researchers to store data on private repositories and to make data available only upon “reasonable request,” or to not share it at all.
@@
A primary barrier that experimentalists in the community face is that the development of control software, and even changing existing software requires significant effort. Specifically, the software tools used to control and measure the experimental devices vary between groups, from software available commercially such as Labber [12], to open-access data acquisition frameworks such as QCoDeS [13], to custom, in-house build packages. While many software companies enable certain kinds of standardized measurements, routines, and data processing, the resulting data file structures are not consistent between groups. Even labeling the various gates in a device (e.g., plungers vs. barriers) is at present a local convention and may lead to confusion when different groups try to use the same data for testing. Any centralized scheme should consider that a data-sharing tool should be almost automatic for researchers to consider transitioning from their currently custom and tailored data acquisition and storage tools to community standards; only this can ensure widespread adoption. One solution that could be implemented is to embed seamless and transparent data sharing within the most commonly used software frameworks (QCoDeS, Labber), and to provide a module easy to integrate into others (for example, in Python and/or Matlab).
@@
Given the above discussions, several core principles should be encouraged for standardizing data. In addition to the high-level FAIR (findable, accessible, interoperable, and reusable) data principles [14, 15], there are several additional requirements specific to experimental data:
@@
    NIST could work with the community to define the ontology and core guidelines under the FAIR principles, along with a living process for improving and revising these guidelines, to enable data sharing across the community.
@@ II.3 Performance metrics
A few simple metrics, such as state identification accuracy (the probability of a classifier identifying the right device topology) and tuning success (the probability of an optimizer or other navigation algorithm getting to the right region of parameter space), have been used for the middle to later stages of tuning [16, 17, 18, 19, 20, 21, 22]. The performance of fine-tuning and gate-tuning processes is often assessed through the final fidelity of the qubit through, e.g., randomized benchmarking. Additional metrics for single-qubit operations could be based on other experimental measurements, such as the number of visible Rabi oscillations, where more observed oscillations indicate superior performance. Similar metrics, e.g., the quality of exchange oscillations, should also be developed for two-qubit gates. However, defining more rigorous quantitative and qualitative metrics and processes to assess the expected overall performance of devices remains an open problem in the field that will require the engagement of the whole community. Focusing on both single-qubit and two-qubit gate standardization will be necessary to avoid calibrating an array of single qubits that can not talk to each other. Once standardization is in place, more structured and algorithmic exploration and improvement of the tuning processes can be achieved.
@@
To date, there have been two approaches to testing the performance of tuning algorithms: (i) direct deployment on experimental devices [23, 18, 19, 24] and (ii) using simulated device data [25, 26] for development and initial assessment of tuning methods followed by testing on experimentally acquired data or directly on experimental devices [16, 27, 17, 21, 22, 20, 28]. The benefit of using simulated data is access to information about what the data represents (e.g., the exact charge occupation or the state of the device) as well as the ability to facilitate a controlled study of how the various changes to the device design or the types, combinations, and prevalence of noise impact the functioning of the tuning algorithms. At the same time, simulation tools are not equipped to cover all aspects of the tuning process and may not capture some of the realistic experimental imperfections [29, 30].
@@
While simulated data might be a good starting point for developing and testing new tuning methods, for benchmarking combining synthetic and experimental datasets might be a more appropriate approach. However, unlike simulated data, experimental data does not come with labels, and manual data labeling is not an easy task. It typically involves a team of annotators, ideally, experts in the subject matter, who review each data point and assign the appropriate label based on labeling guidelines and their understanding of the data. The process can thus be very time- and labor-expensive and result in sub-optimal labels due to the complexity of the quantum dot data. Even the seemingly straightforward task of determining whether a given charge stability diagram represents a good double dot can be challenging as the quality good might have very different meanings for different domain experts. The development of a general, systematic, unbiased, and preferably automated labeling procedure might be necessary if experimental data is to be included in a database intended for benchmarking [31]. Such effort needs to be carried out as a collaboration between computational, theoretical, and experimental groups to ensure a satisfying performance of the resulting labeling software.
@@ II.4 Testbeds and collaborations
The virtual testbed approach could complement existing programs such as the Quantum Foundry program [32] and the LPS Qubit Collaboratory [33]. We wish to emphasize that a virtual testbed is only as good as its access to and validation with experimental systems that are representative of devices of interest. At the same time, the ability to make these virtual systems widely accessible is clear, since software is cheaper than hardware.
@@ II.5 Governance and licensing
When choosing the data-sharing platform it is important to ensure that both sides, i.e., those who contribute and those who access the database, have access to the data, making it very easy for everyone to share and reuse the data. An excellent example of a digital ecosystem of openly available data and supporting tools is the Community Resource for Innovation in Polymer Technology (CRIPT) platform – a partnership between the University of Chicago, NIST, Massachusetts Institute of Technology, Dow Inc., and Citrine Informatics [34]. CRIPT provides researchers working in polymer science and engineering with easy access to a large database of polymer data as well as a set of tools for interacting with the CRIPT platform. The European Organization for Nuclear Research (CERN) Open Data Portal [35, 36] and HEPdata [37, 38] which gives access to data produced through the research performed at CERN are two examples of data-sharing platforms from high energy physics. The Phase Field Community Hub [39] provides a space where the phase field practitioners and code developers can share code and compare code output data using a standard set of metrics. These platforms could serve as examples of successfully deployed data- and code-sharing environments.

### v3

# Data Needs and Challenges for Quantum Dot Devices Automation

@@ Abstract
Gate-defined quantum dots are a promising candidate system for realizing scalable, coupled qubit systems and serving as a fundamental building block for quantum computers. However, present-day quantum dot devices suffer from imperfections that must be accounted for, which hinders the characterization, tuning, and operation process. Moreover, with an increasing number of quantum dot qubits, the relevant parameter space grows sufficiently to make heuristic control infeasible. Thus, it is imperative that reliable and scalable autonomous tuning approaches are developed. This meeting report outlines current challenges in automating quantum dot device tuning and operation with a particular focus on datasets, benchmarking, and standardization. We also present insights and ideas put forward by the quantum dot community on how to overcome them. We aim to provide guidance and inspiration to researchers invested in automation efforts.
@@ I Background and motivation
Gate-defined semiconductor quantum dots are a candidate system to realize scalable, coupled qubit systems that serve as a fundamental building block for quantum computers 1, 2. Their potential for leveraging the semiconductor industry’s materials science and fabrication techniques, while promising, remains hard to realize at scale 3. Specifically, current quantum dot devices suffer from variations in the device dimensions and properties, as well as defects. The corresponding disorder in the electronic landscape throughout a device can be overcome by using local gate voltages to tune the device into an operation regime for quantum computing. However, even tuning a double quantum dot device constitutes a nontrivial task, with each dot typically being controlled by at least three metallic gates, each of which influences the number of electrons in the dot, the tunnel coupling to the outer leads, and the interdot tunnel coupling, which are critical parameters for qubit operation.
@@
The current practice of characterizing and tuning quantum dots for qubit operation either manually or using script-based methods is a relatively time-consuming procedure that is inherently impractical for scaling up and applications. Moreover, with an increasing number of quantum dot qubits, the relevant parameter space grows sufficiently to make heuristic control unfeasible. To overcome the limitations of human-driven experimental control, researchers working with semiconducting quantum dot devices have put considerable effort into automating device control and characterization 4.
@@
To date, several automated tuning algorithms for single- and double-quantum-dot devices have been proposed and demonstrated. While initially all efforts focused on developing in-house script-based algorithms that were tailored to a particular device, more recently members of the community have begun to take advantage of the data analysis tools provided by the field of artificial intelligence and, more specifically, supervised and unsupervised machine learning 4. When provided with proper training data, machine-learning-driven and machine-learning-enhanced methods have the flexibility of being applicable to various devices with minimal to no adjustments or retraining 5. Moreover, by learning the governing rules and dynamics directly from the data, such algorithms may be less susceptible to programming errors.
@@
However, machine learning models typically require large, labeled datasets for training, validation, and benchmarking and often lack information about the reliability of the output. Moreover, since the application of machine learning to quantum dot tuning, characterization, and control is a relatively new field of research, it lacks standardized measures of success. The reported success rates vary significantly in both the level and meaning of the reported performance statistics, making it hard (if not impossible) to benchmark the proposed techniques against more traditional tuning approaches or against one another 4. The time is thus ripe to discuss the broadly defined needs and potential next steps in the field of quantum dot device automation that would enable substantive progress.
@@
When it comes to datasets, most characterization and tuning efforts undertaken to date rely on datasets that consist of either simulated data (which may lack important features representing real-world noise and imperfections) or manually labeled experimental data (which might be subject to qualitative and/or erroneous classification). Moreover, with a few exceptions, experimental datasets are not made publicly available. At the same time, systematic benchmarking of tuning methods on standardized datasets, analogous to the MNIST 6 or CIFAR 7 datasets, is a crucial next step on the path to developing reliable and scalable autotuning algorithms for quantum dot devices.
The goal of the Workshop on Advances in Automation of Quantum Dot Devices Control was to serve as a starting point for discussions about the community’s needs and interests 8. For two days stakeholders from industry, academia, and the government interested in research and development of semiconductor quantum computing technologies discussed methods of collaboration and future roadmap development of methods for tuning large-scale devices. Topics discussed during the meeting included:
@@
Figure 1: Workshop thematic map. The workshop discussions revolved around five main themes: science and policy challenges; guidelines for sharing data; developing performance metrics; establishment of testbed and collaborations; and governance and licensing. These themes are connected to allow the flow of data and information. The end result is an infrastructure for facilitating the development, benchmarking, and standardization of the quantum dot automation methods, where the science is the driver (top row), the testbeds and collaborations are critical (bottom row), and the governance (left box) and the guidelines (right box) take their inputs from the science and provide useful help to the testbed and collaborations. For example, theory, computation, and experiment provide a description of needs that can be met through testbeds and collaboration, the latter two of which are facilitated by governance and guidelines.
The workshop discussions, guided by a set of questions first put forward in a Federal Register Notice published on April 13, 2023 9, revolved around five main themes, which we depict in Fig. 1 present in the following sections. In terms of datasets and benchmarking, the workshop participants were asked to identify public or restricted-use datasets related to the various phases of tuning semiconductor quantum dot devices that are available for training and benchmarking new artificial intelligence models or to test hypotheses using data mining/machine learning methods. They were asked to consider the work researchers need to do to access, and then explore the quality of, existing datasets before conducting research with them and to identify what aspects of this work could be reduced or conducted just once so that future researchers can reduce the time needed to complete a research project. The workshop participants were also asked to consider best practices for creating new datasets or linking existing datasets and sharing them with researchers while adhering to local, State, and Federal laws, as well as barriers and limitations that currently exist to data sharing. Finally, they were asked to think about what role the National Institute of Standards and Technology (NIST) can play in developing infrastructure that supports the use of large-scale datasets for research on tuning quantum dot devices.
@@
In addition to the overall comprehensive summary, each thematic section of this report concludes with a set of core recommendations for possible next steps for advancing automation of the quantum dot tuning and characterization. These recommendations do not represent an endorsement by or an official position of NIST or any other agency participating in the workshop. Rather, they should be understood as summarizing the plurality of views from the workshop.
@@
      NIST could work with the community to define the ontology and core guidelines under the FAIR (findable, accessible, interoperable, and reusable) principles, along with a living process for improving and revising these guidelines, to enable data sharing across the community.
@@ II.1 Science and process challenges
At a high level, quantum dot systems have a variety of engineering and science challenges. However, the largest, pressing challenge is developing a reliable path towards large arrays, in one or two dimensions (or other geometries with scaling at least proportional to the linear size of the system), of quantum dots that are tunable into the quantum coherent, few-electron regime 10.
An example of the challenge ahead was characterized by the difficulties of direct emulation of the Hubbard model. In 2019, Vandersypen and co-authors demonstrated Nagaoka ferromagnetism in a small quantum dot array with four dots 11. This required building a Hamiltonian-level understanding of the device and maintaining the device in a narrow regime of Hamiltonian parameters throughout the experiment. The time to tune up, and retune, the device becomes a potential limiting factor for scaling to larger numbers of qubits (or, in this case, Hubbard model sites).
@@ II.2 Guidelines for sharing data
Advances in the automation of the characterization and tuning of quantum dot devices cannot happen without the joint effort of the whole community. The development of specialized automation techniques, both machine-learning- and non-machine-learning-driven, is at present stymied by the lack of high-quality relevant datasets. Although some groups already share their data using general-purpose open repositories such as Zenodo 12 or Open Science Framework 13, such data is typically unlabeled and limited in scope to include only the good data that was measured to create the final version of the paper. While a lot of things can go wrong with measurement, the negative counter-examples representing bad data (e.g., poor quality measurement, measurements over incorrectly set parameters) or failure modes are rarely made available. This makes it challenging to learn from mistakes and results in multiple groups reproducing the same errors. Moreover, in some groups, it is still a fairly common practice for researchers to store data on private repositories and to make data available only upon “reasonable request,” or to not share it at all.
@@
A primary barrier that experimentalists in the community face is that the development of control software, and even changing existing software requires significant effort. Specifically, the software tools used to control and measure the experimental devices vary between groups, from software available commercially such as Labber 14, to open-access data acquisition frameworks such as QCoDeS 15, to custom, in-house build packages. While many software companies enable certain kinds of standardized measurements, routines, and data processing, the resulting data file structures are not consistent between groups. Even labeling the various gates in a device (e.g., plungers vs. barriers) is at present a local convention and may lead to confusion when different groups try to use the same data for testing. Any centralized scheme should consider that a data-sharing tool should be almost automatic for researchers to consider transitioning from their currently custom and tailored data acquisition and storage tools to community standards; only this can ensure widespread adoption. One solution that could be implemented is to embed seamless and transparent data sharing within the most commonly used sofftware frameworks (QCoDeS, Labber), and to provide a module easy to integrate into others (for example, in Python and/or Matlab).
@@
Given the above discussions, several core principles should be encouraged for standardizing data. In addition to the high-level FAIR data principles 16, 17, there are several additional requirements specific to experimental data:
@@
    NIST could work with the community to define the ontology and core guidelines under the principles of FAIR TASC, along with a living process for improving and revising these guidelines, to enable data sharing across the community.
@@ II.3 Performance metrics
A few simple metrics, such as state identification accuracy (the probability of a classifier identifying the right device topology) and tuning success (the probability of an optimizer or other navigation algorithm getting to the right region of parameter space), have been used for the middle to later stages of tuning 18, 19, 20, 21, 22, 23, 24. The performance of fine-tuning and gate-tuning processes is often assessed through the final fidelity of the qubit through, e.g., randomized benchmarking. Additional metrics for single-qubit operations could be based on other experimental measurements, such as the number of visible Rabi oscillations, where more observed oscillations indicate superior performance. Similar metrics, e.g., the quality of exchange oscillations, should also be developed for two-qubit gates. However, defining more rigorous quantitative and qualitative metrics and processes to assess the expected overall performance of devices remains an open problem in the field that will require the engagement of the whole community. Focusing on both single-qubit and two-qubit gate standardization will be necessary to avoid calibrating an array of single qubits that can not talk to each other. Once standardization is in place, more structured and algorithmic exploration and improvement of the tuning processes can be achieved.
@@
To date, there have been two approaches to testing the performance of tuning algorithms: (i) direct deployment on experimental devices 25, 20, 21, 26 and (ii) using simulated device data 27, 28 for development and initial assessment of tuning methods followed by testing on experimentally acquired data or directly on experimental devices 18, 29, 19, 23, 24, 22, 30. The benefit of using simulated data is access to information about what the data represents (e.g., the exact charge occupation or the state of the device) as well as the ability to facilitate a controlled study of how the various changes to the device design or the types, combinations, and prevalence of noise impact the functioning of the tuning algorithms. At the same time, simulation tools are not equipped to cover all aspects of the tuning process and may not capture some of the realistic experimental imperfections 31, 32.
@@
While simulated data might be a good starting point for developing and testing new tuning methods, for benchmarking combining synthetic and experimental datasets might be a more appropriate approach. However, unlike simulated data, experimental data does not come with labels, and manual data labeling is not an easy task. It typically involves a team of annotators, ideally, experts in the subject matter, who review each data point and assign the appropriate label based on labeling guidelines and their understanding of the data. The process can thus be very time-consuming and labor-intensive, and result in sub-optimal labels due to the complexity of the quantum dot data. Even the seemingly straightforward task of determining whether a given charge stability diagram represents a good double dot can be challenging as the quality good might have very different meanings for different domain experts. Similarly, qubit-specific tuning may be required. For example, tunnel rates between dots are essential for two-qubit gates single-spin qubits 33, 34 and even for individual qubit operation for exchange-only 35 or quantum dot hybrid qubits 36, and the tuning of these tunnel rates interacts with the optimum choice for gate voltage control pulses. An example of this type of interaction is visible in so-called “fingerprint plots” for the exchange-only qubit 37. The development of a general, systematic, unbiased, and preferably automated labeling procedure might be necessary if experimental data is to be included in a database intended for benchmarking 38. Such effort needs to be carried out as a collaboration between computational, theoretical, and experimental groups to ensure a satisfying performance of the resulting labeling software.
@@ II.4 Testbeds and collaborations
The virtual testbed approach could complement existing programs such as the Quantum Foundry program 39 and the LPS Qubit Collaboratory 40. We wish to emphasize that a virtual testbed is only as good as its access to and validation with experimental systems that are representative of devices of interest. At the same time, the ability to make these virtual systems widely accessible is clear, since software is cheaper than hardware.
@@ II.5 Governance and licensing
When choosing the data-sharing platform it is important to ensure that both sides, i.e., those who contribute and those who access the database, have access to the data, making it very easy for everyone to share and reuse the data. An excellent example of a digital ecosystem of openly available data and supporting tools is the Community Resource for Innovation in Polymer Technology (CRIPT) platform – a partnership between the University of Chicago, NIST, Massachusetts Institute of Technology, Dow Inc., and Citrine Informatics 41. CRIPT provides researchers working in polymer science and engineering with easy access to a large database of polymer data as well as a set of tools for interacting with the CRIPT platform. The European Organization for Nuclear Research (CERN) Open Data Portal 42, 43 and HEPdata 44, 45 which gives access to data produced through the research performed at CERN are two examples of data-sharing platforms from high energy physics. The Phase Field Community Hub 46 provides a space where the phase field practitioners and code developers can share code and compare code output data using a standard set of metrics. These platforms could serve as examples of successfully deployed data- and code-sharing environments.